# PySCIPOpt-ML: Embedding Trained Machine Learning Models into Mixed-Integer Programs

### v1

@@ 4 Instance Library
Accompanying PySCIPOpt-ML, we present a new instance library SurrogateLib [^7] https://zenodo.org/records/10357875 consisting of MIPs that contain embedded ML predictors as either constraints or components of the objective function. These instances are inspired by diverse real-world applications, including auto manufacturing (see Section 4.1), water treatment, and adversarial attacks. By avoiding synthetically generated data to train the ML predictors, we aim to showcase the practical benefits of embedding predictive models in optimisation problems.
@@
Alongside the library, we also include instance generators of the introduced problems that allow users to create homogeneous instances of varying difficulty. Thus, PySCIPOpt-ML and SurrogateLib present an opportunity to generate semi-realistic MIPs of controllable complexity by adjusting the size of the embedded ML predictors as well as the parameters of the instances. This way, we provide the MIP community with a set of homogeneous instances that are sufficiently diverse and relatively easy to solve without being trivial. In the following, we will present one of SurrogateLib’s base problems.
4.1 Auto Manufacturer Example
Let us take the point of view of an auto manufacturer that needs to design a new vehicle. Based on historical data about past sales, the manufacturer aims to maximise the number of vehicles sold, while ensuring that the designed vehicle is sufficiently different to any other popular vehicle already on the market. Additionally, the manufacturer wants to ensure that the designed vehicle has a high resell value and is fuel-efficient.
Let $x\in\mathbb{R}^{n}$ be the vector of variable vehicle features that the manufacturer can control. In our case, we have $n=10$ features: vehicle type, engine size, horsepower, wheelbase, width, length, curb weight, fuel capacity, efficiency and power performance factor. Every feature has a lower bound $l_{i}$ and an upper bound $u_{i}$, that is, $l\leq x\leq u$ with $l,u\in\mathbb{R}^{n}$.
Furthermore, let $z_{\text{price}}$, $z_{\text{resell}}$, and $z_{\text{sold}}$ be the features of the vehicle that the manufacturer cannot control. In this example, these uncontrollable features are the price of the vehicle, its resell value, and the number of sold vehicles.
Since the relationship between $x$ and $(z_{\text{price}},z_{\text{resell}},z_{\text{sold}})$ is unknown, the auto manufacturer uses ML to learn from real-world historic data [^8] https://www.kaggle.com/datasets/gagandeep16/car-sales. To do so, a ML predictor, $g$, is trained that takes as input the controllable features of the vehicle and outputs the predicted price. Two more ML predictors, $f$ and $h$, are trained, which take as input the controllable features and the predicted price of the vehicle, and output the predicted amount of vehicles sold and predicted resell value respectively.
Recall that the manufacturer aims to design a vehicle different to what already exists, requiring us to introduce additional constraints. For a finite index set $\mathcal{V}$, we denote by $\bar{x}^{j}\in\mathbb{R}^{10},j\in\mathcal{V}$, the controllable features of the $|\mathcal{V}|$ many most popular vehicles. To ensure that $x$ is sufficiently different to each $\bar{x}^{j}$, we consider the relative difference of each feature and enforce that the cumulative difference among all features are above a given threshold $\gamma_{j}$. Formally, we obtain the constraints
$\sum_{i=1}^{10}\frac{|x_{i}-\bar{x}_{i}^{j}|}{u_{i}-l_{i}}\geq\gamma_{j},\quad\forall j\in\mathcal{V}.$ (1)
Note that even though the absolute value function is non-linear, it can be easily modelled using linear constraints and binary variables.
Finally, let $\alpha$ be the minimum ratio of the resell value compared to the actual purchase price, and let $\beta$ be the minimum level of fuel-efficiency. To ensure that the resulting vehicle has a high enough resell value and is fuel-efficient enough, we need to add the constraints:
$z_{\text{resell}}$ $\geq\alpha z_{\text{price}},$ (2)
$x_{\text{eff}}$ $\geq\beta.$ (3)
Here, $x_{\text{eff}}$ denotes the entry in $x$ that corresponds to the fuel efficiency feature.
The resulting optimisation problem with embedded ML predictors is then given by:
@@ 4.1 Auto Manufacturer Example
s.t. $z_{\text{price}}=g(x),$
$z_{\text{sold}}=f(x,z_{\text{price}}),$
$z_{\text{resell}}=h(x,z_{\text{price}}),$
@@
This parameterised problem can now be used to generate a set of homogeneous instances. By choosing different parameters $\alpha,\beta,\gamma_{j}$ and $\bar{x}_{j}$ for $j\in\mathcal{V}$ as well as different ML models to obtain predictors $f,g$ and $h$, we are able to create instances of varying complexity while still having the same underlying structure.
5 Conclusion
In this paper we have introduced the Python package PySCIPOpt-ML, which helps to automatically embed trained ML predictors in MIPs. We have briefly introduced the API, expanded on some of the core MIP formulations, and walked through a real-world example. In addition, we have introduced a new set of homogeneous semi-realistic MIP instances that form the extendable library SurrogateLib.
Acknowledgements
We thank Mohammed Ghannam, Robet Luce, and Franziska Schlösser for their help with deploying SCIP via Pip. The work for this article has been conducted in the Research Campus MODAL funded by the German Federal Ministry of Education and Research (BMBF) (fund numbers 05M14ZAM, 05M20ZBM).

### v2

@@ 4 Instance Library
Accompanying PySCIPOpt-ML, we present a new instance library, SurrogateLIB  [^5] https://zenodo.org/records/10357875, consisting of MIPs that contain embedded ML predictors as either constraints or components of the objective function.
SurrogateLIB is motivated by the large availability of data in the ML community and the need for more homogeneous non-trivial instance sets in the MIP community. Its goal is to be a library of instances with controllable complexity that is representative of MIPs with embedded ML predictors. To achieve this, semi-realistic MIP scenario generators are constructed in an array of application areas, including auto manufacturing, water treatment, and adversarial attacks, where real-world data is readily available for training the embedded ML predictors. By avoiding synthetically generated data to train the ML predictors, we aim to showcase the practical benefits of embedding predictive models in optimisation problems. An overview of each problem currently contained in SurrogateLIB, including detailed descriptions of the problems and how they are modelled, can be found in Appendix A. The current iteration of SurrogateLIB contains 1016 instances.
@@
Alongside the library, we also include instance generators of the introduced problems that allow users to create homogeneous instances of varying difficulty. All generators feature two random seeds as common arguments. One is for randomising the ML predictor’s training, and the other is for randomising some constraints of the MIP formulation. Thus, PySCIPOpt-ML and SurrogateLIB present an opportunity to generate semi-realistic MIPs of controllable complexity by adjusting the size of the embedded ML predictors as well as the parameters of the instances. This way, we provide a set of homogeneous instances that are sufficiently diverse and relatively easy to solve without being trivial. Table 2 presents a summary for the current generators in SurrogateLIB.
Example Origin Output Type #(Fixed) Input # ML Pred. Brief Description
Adversarial
Attack [42, 43]
MR (0)/256 1
Maximise chance of incorrect
image classification
Auto Manufacturer  [^6] https://www.kaggle.com/datasets/gagandeep16/car-sales R (0)/10 3
Maximise sales of
sufficiently different car
City Planning  [^7] https://www.kaggle.com/datasets/krzysztofjamroz/apartment-prices-in-poland R (7)/14 25
Maximise living quality
of new residents
Function
Approximation [33]
R (0)/5 2
Maximise function while
satisfying equality constraints
Patalable Diet
Problem [44, 22]
R (2)/25 1
Minimise transport cost while meeting
nutritional / palatability requirements
Tree Planting [45] R (6)/7 100 Maximise tree survival rates
Water Potability  [^8] https://github.com/MainakRepositor/Datasets/tree/master C (0)/9 50
Maximise amount of drinkable
water after treatment
Wine Quality [46] R (0)/11 5
Maximise quality of mixed
wine from suppliers
Workload
Dispatching[24]
R (0)/3 48
Maximise minimum efficiency
of compute cores
Table 2: Overview of instance generators available in SurrogateLIB. Output type is labelled as (M)ulti-(R)regression or (C)lassification. #(Fixed) Input refers to the number of MIP variables that are input to each embedded ML predictor, both for those that can take free values and for those that are fixed. #ML Pred. is the number of embedded ML predictors by default in the generator.
5 Computational Experiments
One question that often arises when practitioners deal with surrogate models is the following: How large of a ML predictor can actually be embedded without blowing up solution time? To answer this question, we will present a set of experiments aiming to provide some intuition for the scale of embedded ML predictors that practitioners can expect to optimise over.
To determine the impact of embedding a ML predictor on the solution time, we use SurrogateLIB and progressively embed larger ML predictors until the solution time is 30 minutes or greater. We embed fully connected feed forward neural networks with ReLU activation functions and gradient-boosted decision trees (GBDT). For the neural network, we test both the SOS1 and big-M formulation, where the hidden layer size is fixed to 16. For GBDTs we fix the maximum depth of each estimator to 5. Increasing the size of each ML predictor then corresponds to either introducing an additional hidden layer or an additional estimator. The function approximation problem has been removed so as to leave only semi-real-world examples. GBDT is not used for problems where the embedded ML predictors have fewer than $4$ non-fixed inputs, and for the image based adversarial attack problem.
We fixed the hidden layer sizes of the neural networks and depth of individual estimators in the GDBTs as we wanted only a single axis on which to increase the size of the embedded predictors. Size 16 was chosen as it is a common choice for smaller neural network architectures, and depth 5 was chosen as it is appropriate for a wide range of SurrogateLIB instances. We reduced the experiment to only two types of ML predictors as they admit valid integer linear formulations, can be scaled, and encompass simpler ML predictors, e.g., decision trees.
For all experiments, SCIP 9.0.0 [47] is used for MIP solving with SOPLEX 7.0.0 [47] as the LP solver, PyTorch 2.1.0 [7] is used for training the neural networks, and Scikit-learn 1.4.2 [4] is used for training the GBDTs. All experiments are run on a cluster equipped with Intel Xeon Gold 5122 CPUs with 3.60GHz and 96GB main memory.
During preliminary experiments, we observed a common problem where deeper neural networks would collapse into predicting a constant output. We believe this is caused by vanishing gradients, and is the result of using too deep of a neural network to represent simpler relations. The resulting MIP with the collapsed neural network embedded is often quickly identified as infeasible, as the constant output cannot be changed to satisfy some constraints. Such a result is ultimately misleading for determining the scale of ML predictors that can be embedded. To overcome this issue, we initialised the neural networks with a Xavier uniform distribution [48], introduced gradient clipping [49], and added an L2 penalty. Additionally, we stopped the experiment at a maximum of 15 layers or estimators.
Adversarial Attack Auto Manufacturer City Planning Platatable Diet
ReLU SOS 6 2 0 14*
ReLU Big-M 6 5 0 13*
GBDT - 15+ 2 15+
Tree Planting Water Potability Wine Manufacturer Workload Dispatching
ReLU SOS 0 8 3 0
ReLU Big-M 2* 8 3 0
GBDT - 6 4 -
Table 3: Maximum embedding size for ML predictors in SurrogateLIB such that the resulting MIP can be solved within 30 minutes. Entries with "0" indicate that the starting embedding size could not be solved for the corresponding problem-formulation pair. Entries with "-" indicate that the corresponding experiment was skipped.
*Unreliable results due to collapsing neural network.
The summarised result for the experiment is presented in Table 3. The table shows that the scale of ML predictors that can be practically embedded is highly predictor-dependent. For example, in the auto manufacturer instance, a predictor of scale 2, 5, or 15+ can be solved to optimality in under 30 minutes depending on the embedded predictor type and formulation. Furthermore, we can also observe a dependence on the instance-type. Whereas PySCIPOpt-ML is able to solve instances of the Water Potability problem with relatively large embedded ML predictors, it fails to solve the Workload Dispatching problem with the smallest embedded ML predictor size. This is despite the Water Potability problem having a larger input space for each embedded ML predictor, and more embedded ML predictors. This shows that SurrogateLIB contains a broad range of generators of varying difficulty, making it a versatile tool to test new approaches for solving MIPs with surrogate models.
6 Conclusion
In this paper we have introduced the Python package PySCIPOpt-ML, which automatically embeds trained ML predictors in MIPs. We have briefly introduced the API, expanded on some of the core MIP formulations including novel formulations for ReLU neural networks and the argmax function, and presented a set of computation results that provide an intuition for the scale of ML predictors that can be embedded in practice. For the final point, we introduced a library of homogeneous semi-real-world MIP instance generators, which form the extendable library SurrogateLIB.
Acknowledgements
We thank Mohammed Ghannam, Robert Luce, Julian Manns, and Franziska Schlösser for their help with deploying SCIP via PyPI. The work for this article has been conducted in the Research Campus MODAL funded by the German Federal Ministry of Education and Research (BMBF) (fund numbers 05M14ZAM, 05M20ZBM).
Appendix A SurrogateLIB
A common problem when formulating a MIP formulation of a system is the presence of unknown or highly complex relationships. To address this challenge, ML predictors are often used to approximate these relationships and are then embedded in the MIP as surrogate models. Derived from this terminology, we present SurrogateLIB, a new library of MIP instances.
SurrogateLIB is motivated by the large availability of data in the ML community and the need for more homogeneous non-trivial instance sets in the MIP community. Its goal is to be a library of instances with controllable complexity that is representative of MIPs with embedded ML predictors. To achieve this, semi-realistic MIP scenario generators are constructed in an array of application areas where real-world data is available for training the embedded ML predictors. Each generator features two random seeds as common arguments, which control the ML predictor initialisation and some randomisation of the MIP budget constraints.
The naming convention of an instance from SurrogateLIB requires us to introduce some notation. Let $A$ be the problem type, $B$ be the problem parameters, $C$ be the ML predictor type, $D$ be the predictor parameters, $E$ be the ML framework, and $F$, $G$ be the data and training random seeds. An instance from SurrogateLIB is then named:
$A\_B\_C\_D\_E\_F\_G.mps$
We explicitly list the problem types ($A$) and the associated problem parameters ($B$) in Table 4, where thorough definitions of each generator are given in the subsections of the Appendix.
$A$ adversarial auto city function palatable tree water wine workload
$B$ n - n n - n n n, m -
Table 4: Simplified naming scheme of each problem type and the problem parameters as detailed in the full descriptions
The predictor types ($C$) that we have currently embedded as part of SurrogateLIB are: linear (linear regression), dt (decision tree), gbdt (gradient boosted decision tree), rf (random forest), mlp-sos / mlp-bigm (feed forward neural network with ReLU activation functions and the SOS1 / Big-M formulation respectively). The parameters ($D$) for each predictor that follow in the naming scheme are given in Table 5.
$C$ linear dt gbdt rf mlp-sos mlp-bigm
$D$ depth # estimators, depth # estimators, depth # layers, layer size # layers, layer size
Table 5: ML predictor parameters for SurrogateLIB
The list of ML frameworks, i.e., valid values of $E$, consists of: sk (scikit-learn [4]), torch (PyTorch [7]), keras (Keras [9]), lgb (LightGBM [6], and xgb (XGBoost [5]). The data and training seed, $F$ and $G$, can independently take the values 00 and $1$.
We now present each generator for SurrogateLIB in the following subsections.
A.1 Adversarial Example
An adversarial example is a purposefully altered input instance to a trained ML predictor such that its resulting label is incorrect. Adversarial examples are of extreme importance to many applications [43], with an entire field devoted to robust ML training [50]. They feature as a prominent example for embedded ML predictors in MIPs as they require only the ML predictor and an example input to construct the entire MIP. We use the well-known MNIST dataset [42] in our generator of adversarial example MIPs.
Let $f:[0,1]^{n\times n}\rightarrow\mathbb{R}^{10}$ be the trained ML predictor that takes as input an $n\times n$ image of a black and white photo and outputs the probability of each digit $[0,9]$ being the image, where $n\in\mathbb{N}$. The variables $\bar{x}_{i,j}\in[0,1]$ for $i,j\in[n]$ are the pixel values of the input instance after perturbation, where $x_{i,j}\in[0,1]$ for $i,j\in[n]$ is the given image. The variables $d_{i,j}\in[0,1]$ for $i,j\in[n]$ represent the difference between $\bar{x}_{i,j}$ and $x_{i,j}$, and $\delta\in\mathbb{R}_{\geq 0}$ represents the total difference budget. Finally, $y\in\mathbb{R}^{10}$ are the output probabilities, where without loss of generality $y_{1}$ is the true label and $y_{2}$ is the largest valued incorrect label for the base image. The MIP is defined as:
@@ A.1 Adversarial Example
s.t. $\bar{x}_{i,j}-x_{i,j}\leq d_{i,j},\;\;\forall i,j\in[n]$
$x_{i,j}-\bar{x}_{i,j}\leq d_{i,j},\;\;\forall i,j\in[n]$
$\sum_{i=1}^{n}\sum_{j=1}^{n}d_{i,j}\leq\delta$
$y=f(x).$
A.2 Auto Manufacturer
For this generator, we take the point of view of an auto manufacturer that needs to design a new vehicle. Based on historical data about past sales, the manufacturer aims to maximise the number of vehicles sold, while ensuring that the designed vehicle is sufficiently different to any other popular vehicle already on the market. Additionally, the manufacturer wants to ensure that the designed vehicle has a high resell value and is fuel-efficient.
Let $x\in\mathbb{R}^{n}$ be the vector of variable vehicle features that the manufacturer can control, where for our case $n=10$. These features are: vehicle type, engine size, horsepower, wheelbase, width, length, curb weight, fuel capacity, efficiency and power performance factor. Every feature has a lower bound $l_{i}\in\mathbb{R}$ and an upper bound $u_{i}\in\mathbb{R}$, that is, $l\leq x\leq u$ with $l,u\in\mathbb{R}^{n}$. Furthermore, let $y_{\text{price}}\in\mathbb{R}$, $y_{\text{resell}}\in\mathbb{R}$, and $y_{\text{sold}}\in\mathbb{R}$ be the features of the vehicle that the manufacturer cannot control. In this example, these uncontrollable features are the price of the vehicle, its resell value, and the number of sold vehicles.
Since the relationship between $x$ and $(y_{\text{price}},y_{\text{resell}},y_{\text{sold}})$ is unknown, the auto manufacturer uses ML to learn from real-world historic data [^9] https://www.kaggle.com/datasets/gagandeep16/car-sales. To do so, a ML predictor, $f:\mathbb{R}^{10}\rightarrow\mathbb{R}$, is trained that takes as input the controllable features of the vehicle and outputs the predicted price. Two more ML predictors, $g:\mathbb{R}^{11}\rightarrow\mathbb{R}$ and $h:\mathbb{R}^{11}\rightarrow\mathbb{R}$, are trained, which take as input the controllable features and the predicted price of the vehicle, and output the predicted amount of vehicles sold and predicted resell value respectively.
Recall that the manufacturer aims to design a vehicle different to what already exists, requiring us to introduce additional constraints. For a finite index set $\mathcal{V}$, we denote by $\bar{x}^{j}\in\mathbb{R}^{10},j\in\mathcal{V}$, the controllable features of the $|\mathcal{V}|$ many most popular vehicles. To ensure that $x$ is sufficiently different to each $\bar{x}^{j}$, we consider the relative difference of each feature and enforce that the cumulative difference among all features are above a given threshold $\gamma\in\mathbb{R}$. Formally, we obtain the constraints
$\sum_{i=1}^{10}\frac{|x_{i}-\bar{x}_{i}^{j}|}{u_{i}-l_{i}}\geq\gamma,\quad\forall j\in\mathcal{V}.$ (1)
Note that even though the absolute value function is non-linear, it can be modelled using linear constraints and binary variables. Finally, let $\alpha\in[0,1]$ be the minimum ratio of the resell value compared to the actual purchase price, and let $\beta\in\mathbb{R}$ be the minimum level of fuel-efficiency, where without loss of generality $x_{0}$ is the feature index for fuel-efficiency. The MIP is defined as:
[…]
s.t. $y_{\text{price}}=f(x)$
$y_{\text{sold}}=g(x,y_{\text{price}})$
$y_{\text{resell}}=h(x,y_{\text{price}})$
$y_{\text{resell}}\geq\alpha y_{\text{price}}$
$x_{0}\geq\beta$
@@ A.2 Auto Manufacturer
A.3 City Manager
For this generator, we take the point of view of a city council that has been tasked with building amenities missing in the community. These amenities include two schools, two doctor clinics, two post-offices, two kindergartens, two restaurants, two pharmacies, and a college. As we do not have a direct measure for quality of life improvements by the distance to these amenities, we use the price of apartments as a proxy, for which data is readily available  [^10] https://www.kaggle.com/datasets/krzysztofjamroz/apartment-prices-in-poland.
Let $f:\mathbb{R}^{14}\rightarrow\mathbb{R}$ be the trained ML predictor that takes as input a feature vector of an apartment and outputs the predicted price. A feature vector $x\in\mathbb{R}^{14}$ of an apartment consists of the square meters, number of rooms, floor level, whether the apartment has parking, a balcony, or security, and the shortest distance to the town center, a doctor’s office, a post-office, a kindergarten, a restaurant, a pharmacy, and a college. We construct an imaginary town of $n\in\mathbb{N}$ many apartments in an $m\times m$-km grid, where $m\in\mathbb{R}_{\geq 0}$. The features of each apartment that are not distance based are randomised and fixed. This includes the location of the apartment, and therefore the distance to the town center. The output of the ML predictor for apartment $i\in[n]$ is represented by the variable $y_{i}\in\mathbb{R}$.
[…]
s.t. $y_{i}=f(x^{i}),\;\;\forall i\in[n]$
$\sum_{k=1}^{2}|b^{j}_{k,0}-b^{j}_{k,1}|\geq\gamma,\;\;\forall j\in\mathcal{J}\setminus\{\text{college}\}$
$x^{i}_{j}=\text{min}(\sum_{k=1}^{2}|b^{j}_{k,0}-a_{i,k}|,\sum_{k=1}^{2}|b^{j}_{k,1}-a_{i,k}|)\;\;\forall j\in\mathcal{J}\setminus\{\text{college}\},i\in[n]$
$x^{i}_{\text{college}}=\sum_{k=1}^{2}|b^{\text{college}}_{k,0}-a_{i,k}|\;\;\forall i\in[n]$
Note that even though the absolute value and min functions are non-linear, they can be modelled via linear constraints and additional binary variables.
A.4 Function Approximation
As opposed to all other members of SurrogateLIB, this MIP is entirely contrived from random data. It was the MIP used in [33] to evaluate the impact of different big-M formulations and bound-tightening procedures on solver performance.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be trained ML predictors, where $n\in\mathbb{N}$. Each predictor approximates a different random quadratic function of the form $x^{\intercal}Qx+a^{\intercal}x+b$, where $x\in\mathbb{R}^{n}$, $Q\in\mathbb{R}^{n\times n}$, $a\in\mathbb{R}^{n}$, and $b\in\mathbb{R}$. The values of $Q,a$, and $b$ are samples from a uniform distribution. Additionally, a constant value $c\in\mathbb{R}$ is generated. The MIP is defined as:
[…]
A.5 Palatable Diet
For this generator, we take the view point of the World Food Programme that must decide the quantities of different food types that it will deliver. This scenario phrased as an optimisation problem has been discussed heavily in [44], and was a leading example for embedded ML predictors in MIP for the framework OptiCL [22]. Borrowing data that they have made publicly available, we construct a simplified variant of the original food delivery problem, which was coined The Palatable Diet Problem. The name is derived by extending the traditional diet problem via a minimum palatability constraint.
Let $x\in\mathbb{R}^{25}_{\geq 0}$ be a feature vector that represents a basket of 25 types of food. Additionally, let $n_{i,j}\in\mathbb{R}_{\geq 0}$ be the nutritional value of nutrient $j\in\mathcal{J}$ for one unit of food $i\in[25]$, where for our scenario we consider 12 nutrients, i.e., $|\mathcal{J}|=12$. The value $\gamma_{j}\in\mathbb{R}_{\geq 0}$ is a minimum requirement of nutrient $j\in\mathcal{J}$. While some baskets of food would meet nutritional requirements, they are undesirable from a consumers perspective. Therefore, a palatability requirement is enforced, with this requirement being derived from data due to the complexity of defining such a constraint. Let $f:\mathbb{R}^{25}\rightarrow\mathbb{R}$ be a trained ML predictor that predicts the palatability of a given food basket, and $\beta\in\mathbb{R}$ be some minimum requirement. Additionally, let $c_{i}\in\mathbb{R}_{\geq 0}$ be the transportation cost of a single unit of food for $i\in[25]$. The MIP is defined as:
[…]
$\;\;\;\;\;\;\;f(x)\geq\beta$
$\;\;\;\;\;\;\;x_{\text{salt}}=5$
$\;\;\;\;\;\;\;x_{\text{sugar}}=20$
A.6 Tree Planting
For this generator, we take the point of view of an individual responsible for replanting a set of trees on a stretch of empty land. The goal of this problem is to plant a diverse set of trees, while respecting a budget constraint, and maximising the total amount of surviving trees. Thanks to studies on the effect of soil sterilisation and light levels on tree survival rates [45], we have data readily available for such a problem.
Let $x_{i}\in\mathbb{R}^{7}$ be a feature vector of planting location $i\in[n]$, where $n\in\mathbb{N}$. We randomly generate a grid of planting locations with randomised feature values, where the binary decision of whether the location is sterilised set to false. A sterilisaion budget is then introduced and denoted by $\beta\in\mathbb{N}$. For each planting location we must plant one of four tree species $k\in[4]$. Each species has a minimum expected survival requirement $\gamma_{k}\in\mathbb{R}_{\geq 0}$, and a planting cost $c_{k}\in\mathbb{R}_{\geq 0}$. For each species, a trained ML predictor $f_{k}:\mathbb{R}^{7}\rightarrow\mathbb{R}$ is constructed that predicts the probability that a tree of species $k\in[4]$ survives in a given planting location. To model the MIP we introduce variables $s_{i,k}\in\mathbb{R}$ representing the (adjusted) predicted survival rate of tree species $k$ in location $i$, and $s^{\prime}_{i,k}$ to represent the adjusted survival rates. Binary variables $p_{ik}\in\{0,1\}$ represent whether tree species $k\in[4]$ was planted in location $i\in[n]$. The MIP is defined as:
[…]
$\;\;\;\;\;\;\;p_{i,k}=1\xrightarrow{}s^{\prime}_{i,k}\leq s_{i,k},\;\forall i\in[n],k\in[4]$
$\;\;\;\;\;\;\;p_{i,k}=0\xrightarrow{}s^{\prime}_{i,k}\leq 0,\;\forall i\in[n],k\in[4]$
$\;\;\;\;\;\;\;\sum_{k=1}^{4}c_{k}p_{i,k},\;\forall i\in[n]$
$\;\;\;\;\;\;\;\sum_{k=1}^{4}p_{i,k}=1,\;\forall i\in[n]$
$\;\;\;\;\;\;\;\sum_{i=1}^{n}x^{\text{sterile}}_{i}\leq\beta,$
$\;\;\;\;\;\;\;f_{k}(x_{i})=s_{i,k},\;\forall i\in[n],k\in[4]$
A.7 Water Potability
For this generator, we take the point of view of a health organisation that is treating drinking water. The organisation has access to processes that decrease or increase levels of attributes for a given sample of water. There are however limitations on the total amount that each attribute can be increased or decreased. The goal of the optimisation problem is to maximise the amount of drinkable water given a set of non-drinkable water. Data for this problem is publicly available  [^11] https://github.com/MainakRepositor/Datasets/tree/master.
Let $x_{i}\in\mathbb{R}^{9}$ be the feature vector of given water sample, where we have taken $n\in\mathbb{N}$ many undrinkable water samples. For each water sample $i\in[n]$ and feature $j\in[9]$, let $a_{i,j}\in\mathbb{R}_{\geq 0}$ be the positive change after applying treatment, and $b_{i,j}\in\mathbb{R}_{\geq 0}$ be the negative change after applying treatment. For each feature $j\in[9]$ there is a positive and negative budget $\gamma^{+}_{j},\gamma^{-}_{j}\in\mathbb{R}_{\geq 0}$. Before treating the water, the undrinkable samples have values $w_{i}\in\mathbb{R}^{9}$ for sample $i\in[n]$. To determine if the water is drinkable, we have access to a trained ML predictor $f:\mathbb{R}^{9}\rightarrow\mathbb{R}$. The ML predictor takes the feature vector of a water sample as input and outputs whether or not it is drinkable. The binary variables that determine if water sample $i\in[n]$ is drinkable are $y_{i}\in\{0,1\}$. The MIP is defined as:
[…]
s.t. $x_{i,j}=w_{i,j}+a_{i,j}-b_{i,j},\;\;\forall i\in[n],j\in[9]$
$\sum_{i=1}^{n}a_{i,j}\leq\gamma^{+}_{j},\;\;\forall j\in[9]$
$\sum_{i=1}^{n}b_{i,j}\leq\gamma^{-}_{j},\;\;\forall j\in[9]$
$y_{i}=f(x_{i}),\;\;\forall i\in[n]$
A.8 Wine Blending
For this generator, we take the point of view of a wine manufacturer. This manufacturer does not grow the grapes themselves, but rather purchases grapes from other growers and blends the result into a new wine. The grapes vendors sell grapes because a wine made from their grapes alone is predicted to be of poor quality. The objective of this problem is to create a bouquet of wines of fixed size with highest average quality. Each vineyard has an associated costs for purchasing a unit of grapes, and the wine manufacturer has a total budget. We assume that readily available data from features of wine [46] can be transferred to features of grapes for this optimisation problem, and that the features of a resulting wine are a linear combination of those in the blend.
Let $x_{i}\in\mathbb{R}^{11}_{\geq 0}$ be the feature vector of a grape (wine) blend for $i\in[n]$, and $w_{j}\in\mathbb{R}^{11}_{\geq 0}$ be the fixed feature vector of a wine (grape) from vendor $j\in[m]$. The features of a grape (wine) are fixed acidity, volatile acidity, citric acid, residual sugar, chlorides, free sulfur, total sulfur, density, pH, sulphates, and alcohol percentage. To determine wine quality, we have access to a trained ML predictor $f:\mathbb{R}^{11}\rightarrow\mathbb{R}$, which predicts the wine quality of the grape blend $i\in[n]$ from its associated feature vector $x_{i}$. The variables $y_{i}\in[0,10]$ for $i\in[n]$ represent the predicted quality, where $\alpha\in[0,1]$ is a minimum required quality.
Each grape vendor $j\in[m]$ has an amount of available grapes $\gamma_{j}\in\mathbb{R}_{\geq 0}$, and an associated cost per unit $c_{j}\in\mathbb{R}_{\geq 0}$. The variables $b_{i,j}\in\mathbb{R}_{\geq 0}$ represent the amount of grapes used for wine blend $i\in[n]$ from vendor $j\in[m]$, and $\beta\in\mathbb{R}_{\geq 0}$ is the total budget for the purchasing of all grapes. We assume that all created wine blends are of the same size. The MIP is defined as:
[…]
s.t. $x^{i}=\sum_{j=1}^{m}b_{i,j}w_{j},\;\;\forall i\in[n]$
$\sum_{i=1}^{n}b_{i,j}\leq\gamma_{j},\;\;\forall j\in[m]$
$\sum_{i=1}^{n}\sum_{j=1}^{m}c_{j}b_{i,j}\leq\beta,$
$\sum_{j=1}^{m}b_{i,j}=1,\;\;\forall i\in[n]$
$y_{i}\geq\alpha,\;\;\forall i\in[n]$
$y_{i}=f(x_{i}),\;\;\forall i\in[n]$
A.9 Workload Dispatching
For this generator, we take the point of view of a job scheduler that must issue compute jobs to individual compute cores. The objective of the optimisation problem is to maximise the minimum efficiency of all cores. Similar to the palatable diet problem for OptiCL [22], the Workload Dispatching Problem served as the leading example for the EML framework [24]. Borrowing data that they have publicly available [^12] https://bitbucket.org/m_lombardi/eml-aij-2015-resources/src/master/training/, we construct instances of the problem.
As efficiency of a compute core is difficult to determine apriori, a ML predictor is used. Let $x_{i}\in\mathbb{R}^{3}$ be the feature vector of a set of jobs assigned to compute core $i\in[n]$ for $n\in\mathbb{N}$. The features are the average clock (equivalently cycles) per instruction (CPI) of jobs assigned to the compute core, the average CPI of neighbouring cores, and the average CPI of non-neighbouring cores. The neighbourhood of a core $i\in[n]$ is defined as $\mathcal{N}(i)\subseteq[n]$. To obtain the feature sets, jobs must first be distributed to each core, where an equal amount of jobs are assigned to each core. For this instance $m\in\mathbb{N}$ many jobs are randomly generated with varying CPIs $c_{j}$ for $j\in[m]$. The variables $x_{i,j}\in\{0,1\}$ represent whether job $j\in[m]$ is assigned to compute core $i\in[n]$. A trained ML predictor $f:\mathbb{R}^{3}\rightarrow\mathbb{R}$ predicts the efficiency $y_{i}\in\mathbb{R}$ of compute core $i\in[n]$ after all cores are assigned jobs. Finally, the variable $e\in\mathbb{R}$ represents the minimum efficiency over all cores. The MIP is defined as:
[…]
s.t. $\sum_{i=1}^{n}x_{i,j}=1,\;\;\forall j\in[m]$
$\sum_{j=1}^{m}x_{i,j}=\frac{m}{n},\;\;\forall i\in[n]$
$x_{i}^{\text{avg}}=\frac{n}{m}\sum_{j=1}^{m}c_{j}x_{i,j},\;\;\forall i\in[n]$
$x_{i}^{\text{avg-neigh}}=\frac{1}{|\mathcal{N}(i)|}\sum_{i^{\prime}\in\mathcal{N}(i)}x_{i^{\prime}}^{\text{avg}},\;\;\forall i\in[n]$
$x_{i}^{\text{avg-far}}=\frac{1}{n-|\mathcal{N}(i)|}\sum_{i^{\prime}\in[n]\setminus\mathcal{N}(i)}x_{i^{\prime}}^{\text{avg}},\;\;\forall i\in[n]$
$y_{i}=f(x_{i}^{\text{avg}},x_{i}^{\text{avg-neigh}},x_{i}^{\text{avg-far}})$
$e\leq y_{i},\;\;\forall i\in[n]$